# Advancing Post-OCR Correction: A Comparative Study of Synthetic Data

### v1

@@ 1 Introduction
Crowdsourcing has been the primary approach to acquire post-OCR training data in this domain Clematide et al. (2016); Richter et al. (2018); Maheshwari et al. (2022). While this can provide highly accurate training data, the process is often time consuming and expensive. With the advent of Transformer architecture and attention mechanisms Vaswani et al. (2017), deep learning models have emerged as the standard approach for post-OCR tasks. These models require large amounts of data for training. Thus, synthetic data has been increasingly adopted in this context. D’hondt et al. (2017); davydkin2023data; Jasonarson et al. (2023). However, most existing literature on generating synthetic data relies on additional existing data for generation, and no comprehensive comparison has been made between different methods to understand how the synthetic data generated in various ways affects post-OCR performance.
@@ 2 Related Work
For the task of text denoising, the primary method for constructing synthetic data is noise injection Izumi et al. (2003). This technique involves inserting errors into clean text to generate training data pairs. D’hondt et al. (2017) added artificial OCR errors into sentences using a random process, while Jasonarson et al. (2023) focused on the low-resource Icelandic language for post-OCR tasks. The authors extracted OCR errors from real digitised documents and inserting them into clean text in similar proportions. Grundkiewicz et al. (2019) analyzed real data to create replacement sets for each word. davydkin2023data adopted a different approach, developing a system to generate handwritten image data which were then processed with OCR. The resulting OCR outputs and the original texts were subsequently used to train a T5 model Raffel et al. (2020) for correction purposes. Ignat et al. (2022) synthesized text image datasets by manipulating parameters, like font spacing and image saturation, then compare the original text with the OCR output text, to evaluate OCR systems’ performance on various low-resource languages, and they enhanced Machine Translation (MT) through backtranslation Sennrich et al. (2015).

### v2

@@ 1 Introduction
Crowdsourcing has been the primary approach to acquire post-OCR training data in this domain Clematide et al. (2016); Richter et al. (2018); Maheshwari et al. (2022). While this can provide highly accurate training data, the process is often time consuming and expensive. With the advent of Transformer architecture and attention mechanisms Vaswani et al. (2017), deep learning models have emerged as the standard approach for post-OCR tasks. These models require large amounts of data for training. Thus, synthetic data has been increasingly adopted in this context. D’hondt et al. (2017); Davydkin et al. (2023); Jasonarson et al. (2023). However, most existing literature on generating synthetic data relies on additional existing data for generation, and no comprehensive comparison has been made between different methods to understand how the synthetic data generated in various ways affects post-OCR performance.
@@ 2 Related Work
For the task of text denoising, the primary method for constructing synthetic data is noise injection Izumi et al. (2003). This technique involves inserting errors into clean text to generate training data pairs. D’hondt et al. (2017) added artificial OCR errors into sentences using a random process, while Jasonarson et al. (2023) focused on the low-resource Icelandic language for post-OCR tasks. The authors extracted OCR errors from real digitised documents and inserting them into clean text in similar proportions. Grundkiewicz et al. (2019) analyzed real data to create replacement sets for each word. Davydkin et al. (2023) adopted a different approach, developing a system to generate handwritten image data which were then processed with OCR. The resulting OCR outputs and the original texts were subsequently used to train a T5 model Raffel et al. (2020) for correction purposes. Ignat et al. (2022) synthesized text image datasets by manipulating parameters, like font spacing and image saturation, then compare the original text with the OCR output text, to evaluate OCR systems’ performance on various low-resource languages, and they enhanced Machine Translation (MT) through backtranslation Sennrich et al. (2015).